# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 5.3 Estimate for the supports of densities
Recall $A=m^{-\frac{2}{3\gamma-4}}$, $A^{\gamma-1}m^{2}=m^{\frac{4\gamma-6}{3\gamma-4}}$. Thus when $m$ is small enough, for almost every $y\in\Omega_{m}^{B}:=\left\{y\in\mathbb{R}^{3}\mid{B}y\in\Omega_{m}\right\}$, we have ${\widetilde{\rho_{m}}}^{\gamma-1}(y)\leq\left[V_{\widetilde{\rho_{m}}}(y)+\frac{\kappa_{1}+\epsilon}{2}\right]_{+}$. Now, let $R_{0}$ from Theorem 3.6 (v) bound the support radius of non-rotating minimizer $\sigma$, and choose $\tilde{R}(J)\geq R_{0}$ large enough so that $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$. Using Theorem 5.9, we know up to translation, $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{1}\left(\mathbb{R}^{3}\right)$. In particular, $\forall\delta>0$, when $m$ is small enough, all but mass $\delta$ of $\widetilde{\rho_{m}}$ is forced into a ball of radius $R_{0}$, which is, without loss of generality, centered on 0. Then $\widetilde{\rho_{m}}=\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$, $V_{\widetilde{\rho_{m}}}=V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$. Given $y$ outside the larger ball of radius $\tilde{R}(J)$ centered on 0, $V_{\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}}(y)\leq-\frac{\kappa_{1}+\epsilon}{6}$ due to $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$. For $V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}(y)$, by Theorem 5.14 we can show $\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$ with mass $\delta$. And we pick a $\delta$ satisfying $k\delta^{\frac{2}{3}}\tilde{C}^{\frac{1}{3}}<-\frac{\kappa_{1}+\epsilon}{6}$, where $\tilde{C}$ is the uniform bound of $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}}$ in Theorem 5.14. Thus, when $m$ is small enough, by Proposition 5.20, we have (note $V_{\widetilde{\rho_{m}}}$ is continuous due to Theorem 5.14)
@@
Remark 5.23.
In Lemma 5.18 and Theorem 5.21 we consider the the polytropic law (1) $P(\rho)=K\rho^{\gamma}$. McCann actually considered more general form of pressure $P(\rho)$ [46, Lemma 6.5 and Theorem 6.6]. In [46], $\kappa_{1}$ is replaced by $e_{0}^{\prime}(1^{-})$, where $e_{0}^{\prime}(1^{-})$ is negative and is a upper bound of Lagrange multipliers corresponding to non-rotating minimizers with fixed mass $1$, even in cases where multiple non-rotating minimizers exist (see Theorem 3.6 and Remark 3.9). Nevertheless, since we are dealing with scaling density here, we focus on this polytropic law.
@@
There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\rho_{m}$ is contained in a ball of radius $B\cdot R(J)$, where $B$ is defined in Definition 5.1. In particular, spt $\rho_{m}$ is contained in a ball of a radius which goes to 0 with rate $B$ since $\lim\limits_{m\rightarrow 0}B=0$ if $\gamma>2$.
@@ 6 Existence for Star-Planet Systems
So far, we have dealt with the uniform bound of the supports of density functions with different mass. Our next step is to show the centers of mass can be chosen to be not too far away from the centers of $\Omega_{1-m}$ and $\Omega_{m}$. Thanks to those results, we then can show the constrained minimizers are actually $W^{\infty}$ local energy minimizers, which represent solutions to the reduced Euler-Poisson system (EP’) up to translation due to Theorem 3.1.
@@ 6.1 Estimate for the Center of Mass Separation
In this subsection, we estimate for the centers of mass separation by modifying the arguments in [46, Lemma 6.7 and Proposition 6.8] such that they can be applied in our case.
@@
One can observe when $\epsilon$ is small enough, the function $g_{\epsilon}$ and $g_{\epsilon}^{\prime}$ are analytic and uniformly bounded on $\left\{z\in\mathbb{C}\mid|z|>\frac{1}{4}\right\}$. The result can be check by direct computation and estimation. ∎
@@
Fix $J>0$, let $0<\zeta<\frac{1}{2},\exists\delta>0$, such that $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then $1-\zeta<\frac{\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|}{\left|y_{m}-y_{1-m}\right|}<1+\zeta$, where $y_{m}$ and $y_{1-m}$ are the centers of $\Omega_{m}$ and $\Omega_{1-m},\left|y_{m}-y_{1-m}\right|=\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
@@
Take $m$ small enough such that Corollary 5.24 and Corollary 5.26 provide bound $R(J)$ for the supports of $\rho_{m}$ and $\rho_{1-m}$, and $R:=2R(J)<\frac{\eta}{4}$. Due to Corollary 5.24, we know spt $\rho_{m}\subset B_{R(J)}\left(z_{m}\right)$ for some $z_{m}\in\mathbb{R}^{3}$, then $\bar{x}\left(\rho_{m}\right)\in B_{R(J)}\left(z_{m}\right)$. If $x$ is outside $B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$, then $\left|x-z_{m}\right|\geq\left|x-\bar{x}\left(\rho_{m}\right)\right|-\mid z_{m}$- $\bar{x}\left(\rho_{m}\right)\mid>R(J),\rho_{m}(x)=0$. Therefore, spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$. Similarly, spt $\rho_{1-m}\subset$ $B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$. Now we translate $\rho_{m}$ and $\rho_{1-m}$ separately so that their centers of mass are positioned at $y_{m}$ and $y_{1-m}$ respectively, and denote the shifted functions as $\kappa_{m}$ and $\kappa_{1-m}$. Therefore, $\kappa(m)\coloneq\kappa_{m}+\kappa_{1-m}\in W_{m}$. Due to the proof of Lemma 6.3 below we know $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=c$. Hence the distance between them equals the distance between their projections onto $x_{1}x_{2}$ plane. That is, $d:=\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|=r\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)$, where $r$ is given in Definition 2.1. Since spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$, we have (recall $\mu=m(1-m)$)
$\frac{\mu}{d+2R}\leq G\left(\rho_{m},\rho_{1-m}\right)=\iint\frac{\rho_{m}(x)\cdot\rho_{1-m}(y)}{|x-y|}d^{3}xd^{3}y\leq\frac{\mu}{d-2R}$
@@
$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{1-m}(x)d^{3}x+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{m}(x)d^{3}x$
@@
$\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)=\mu d^{2}\leq I(\rho(m))\leq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+(1-m)R^{2}+mR^{2}=\mu d^{2}+R^{2}$
@@
$\frac{J^{2}}{2\left(\mu d^{2}+R^{2}\right)}\leq T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}\leq\frac{J^{2}}{2\mu d^{2}}$
Similar inequalities hold true for $G\left(\kappa_{m},\kappa_{1-m}\right)$ and $T_{J}(\kappa(m))$, but we replace $d$ by $\eta$, since $\left|\bar{x}\left(\kappa_{m}\right)-\bar{x}\left(\kappa_{1-m}\right)\right|=r\left(\bar{x}\left(\kappa_{1-m}\right)-\bar{x}\left(\kappa_{m}\right)\right)=\left|y_{m}-y_{1-m}\right|=\eta$. Note $\rho(m)$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, we have $E_{J}(\kappa(m))-E_{J}(\rho(m))=$ $G\left(\kappa_{m},\kappa_{1-m}\right)+T_{J}(\kappa(m))-G\left(\rho_{m},\rho_{1-m}\right)+T_{J}(\rho(m))\geq 0$, just as similar as (40). Using the previous estimates and $J^{2}=\mu^{2}\eta$, we have
$\frac{1}{d-2R}+\frac{\mu\eta}{2\left(\mu d^{2}+R^{2}\right)}\leq-\frac{1}{\eta+2R}+\frac{1}{2\eta}$
@@
$\rho(m)\in W_{m}$ implies $\bar{x}\left(\rho_{m}\right)\in\Omega_{m},\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m}$, thus $\frac{1}{2}\leq x\leq\frac{3}{2}$ by the construction of $\Omega_{m}$ and $\Omega_{1-m}$. Let $g_{0}(z)=-\frac{1}{z}+\frac{1}{2z^{2}}+\frac{1}{2}$, by Lemma 6.1 we know $g_{\epsilon}(z)$ converges uniformly to $g_{0}(z)$ as $\epsilon\rightarrow 0$ on $|z|\geq\frac{1}{2}$. One can see $g_{0}(1)=0$, and $\forall\zeta>0$, $\exists C>0$, such that on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$, we have $g_{0}(x)>C$. Due to Lemma 6.1, we have $g_{\epsilon}(x)>\frac{C}{2}>0$ on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$ for $\epsilon$ small enough (i.e. for $m$ small enough since $\epsilon=\frac{R}{\eta}=\frac{Rm^{2}(1-m)^{2}}{J^{2}}$). Thus (3.3.1) implies $x=\frac{d}{\eta}$ has to satisfies $1-\zeta<\frac{d}{\eta}<1+\zeta$, which proves the proposition. ∎

### v2

@@ 5.3 Estimate for the supports of densities
Recall $A=m^{-\frac{2}{3\gamma-4}}$, $A^{\gamma-1}m^{2}=m^{\frac{4\gamma-6}{3\gamma-4}}$. Thus when $m$ is small enough, for almost every $y\in\Omega_{m}^{B}:=\left\{y\in\mathbb{R}^{3}\mid{B}y\in\Omega_{m}\right\}$, we have ${\widetilde{\rho_{m}}}^{\gamma-1}(y)\leq\left[V_{\widetilde{\rho_{m}}}(y)+\frac{\kappa_{1}+\epsilon}{2}\right]_{+}$. Now, let $R_{0}$ from Theorem 3.6 (v) bound the support radius of non-rotating minimizer $\sigma$, and choose $\tilde{R}(J)\geq R_{0}$ large enough so that $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$. Using Theorem 5.9, we know up to translation, $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{1}\left(\mathbb{R}^{3}\right)$. In particular, $\forall\delta>0$, when $m$ is small enough, all but mass $\delta$ of $\widetilde{\rho_{m}}$ is forced into a ball of radius $R_{0}$, which is, without loss of generality, centered on 0. Then $\widetilde{\rho_{m}}=\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$, $V_{\widetilde{\rho_{m}}}=V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$. Given $y$ outside the larger ball of radius $\tilde{R}(J)$ centered on 0, due to $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$, we have
$V_{\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}}(y)\leq-\frac{\kappa_{1}+\epsilon}{6}$
For $V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}(y)$, by Theorem 5.14 we can show $\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$ with mass $\delta$. And we pick a $\delta$ satisfying $k\delta^{\frac{2}{3}}\tilde{C}^{\frac{1}{3}}<-\frac{\kappa_{1}+\epsilon}{6}$, where $\tilde{C}$ is the uniform bound of $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}}$ in Theorem 5.14. Thus, when $m$ is small enough, by Proposition 5.20, we have (note $V_{\widetilde{\rho_{m}}}$ is continuous due to Theorem 5.14)
@@
There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\rho_{m}$ is contained in a ball of radius $B\cdot R(J)$, where $B$ is defined in Definition 5.1. In particular, if $\gamma>2$, spt $\rho_{m}$ is contained in a ball whose radius goes to 0 with rate $B$ since $\lim\limits_{m\rightarrow 0}B=0$.
@@ 6 Existence for Star-Planet Systems
So far, we have dealt with the uniform bound of the supports of density functions with different mass. Our next step is to show the centers of mass can be chosen to be not too far away from the centers of $\Omega_{1-m}$ and $\Omega_{m}$. Thanks to those results, we can then show the constrained minimizers are actually $W^{\infty}$ local energy minimizers, which represent solutions to the reduced Euler-Poisson system (EP’) up to translation due to Theorem 3.1.
@@ 6.1 Estimate for the Center of Mass Separation
In this subsection, we estimate the centers of mass separation by modifying the arguments in [24, Lemma 6.7 and Proposition 6.8] such that they can be applied in our case.
@@
One can observe when $\epsilon$ is small enough, the functions $g_{\epsilon}$ and $g_{\epsilon}^{\prime}$ are analytic and uniformly bounded on $\left\{z\in\mathbb{C}\mid|z|>\frac{1}{4}\right\}$. The result can be checked by direct computation and estimation. ∎
@@
Fix $J>0$, let $0<\zeta<\frac{1}{2},\exists\delta>0$, such that $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then $1-\zeta<\frac{\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|}{\left|y_{m}-y_{1-m}\right|}<1+\zeta$, where $y_{m}$ and $y_{1-m}$ are the centers of $\Omega_{m}$ and $\Omega_{1-m}$, $\left|y_{m}-y_{1-m}\right|=\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
@@
Take $m$ small enough such that Corollary 5.23 and Corollary 5.25 provide bound $R(J)$ for the supports of $\rho_{m}$ and $\rho_{1-m}$, and $R:=2R(J)<\frac{\eta}{4}$. Due to Corollary 5.23, we know spt $\rho_{m}\subset B_{R(J)}\left(z_{m}\right)$ for some $z_{m}\in\mathbb{R}^{3}$, then $\bar{x}\left(\rho_{m}\right)\in B_{R(J)}\left(z_{m}\right)$. If $x$ is outside $B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$, then $\left|x-z_{m}\right|\geq\left|x-\bar{x}\left(\rho_{m}\right)\right|-\mid z_{m}$- $\bar{x}\left(\rho_{m}\right)\mid>R(J)$, $\rho_{m}(x)=0$. Therefore, spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$. Similarly, spt $\rho_{1-m}\subset$ $B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$. Now we translate $\rho_{m}$ and $\rho_{1-m}$ separately so that their centers of mass are positioned at $y_{m}$ and $y_{1-m}$ respectively, and denote the shifted functions as $\kappa_{m}$ and $\kappa_{1-m}$. Therefore, $\kappa(m)\coloneq\kappa_{m}+\kappa_{1-m}\in W_{m}$. Due to the proof of Lemma 6.3 below we know $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=c$. Hence the distance between them equals the distance between their projections onto $x_{1}x_{2}$ plane. That is, $d:=\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|=r\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)$, where $r$ is given in Definition 2.1. Since spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$, we have (recall $\mu_{r}=m(1-m)$)
$\frac{\mu_{r}}{d+2R}\leq G\left(\rho_{m},\rho_{1-m}\right)=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho_{m}(x)\cdot\rho_{1-m}(y)}{|x-y|}\,dx\,dy\leq\frac{\mu_{r}}{d-2R}$
@@
$I(\rho(m))=\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+\int_{\mathbb{R}^{3}}r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{1-m}(x)\,dx+\int_{\mathbb{R}^{3}}r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{m}(x)\,dx$
@@
$\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)=\mu_{r}d^{2}\leq I(\rho(m))\leq\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+(1-m)R^{2}+mR^{2}=\mu_{r}d^{2}+R^{2}$
@@
$\frac{J^{2}}{2\left(\mu_{r}d^{2}+R^{2}\right)}\leq T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}\leq\frac{J^{2}}{2\mu_{r}d^{2}}$
Similar inequalities hold true for $G\left(\kappa_{m},\kappa_{1-m}\right)$ and $T_{J}(\kappa(m))$, but we replace $d$ by $\eta$, since $\left|\bar{x}\left(\kappa_{m}\right)-\bar{x}\left(\kappa_{1-m}\right)\right|=r\left(\bar{x}\left(\kappa_{1-m}\right)-\bar{x}\left(\kappa_{m}\right)\right)=\left|y_{m}-y_{1-m}\right|=\eta$. Note $\rho(m)$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, based on the energy decomposition similar to (41), we have
$E_{J}(\kappa(m))-E_{J}(\rho(m))=-G\left(\kappa_{m},\kappa_{1-m}\right)+T_{J}(\kappa(m))-G\left(\rho_{m},\rho_{1-m}\right)+T_{J}(\rho(m))\geq 0$
Using the previous estimates and $J^{2}=\mu_{r}^{2}\eta$, we have
$\frac{1}{d-2R}+\frac{\mu_{r}\eta}{2\left(\mu_{r}d^{2}+R^{2}\right)}\leq-\frac{1}{\eta+2R}+\frac{1}{2\eta}$
@@
$\rho(m)\in W_{m}$ implies $\bar{x}\left(\rho_{m}\right)\in\Omega_{m},\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m}$, thus $\frac{1}{2}\leq x\leq\frac{3}{2}$ by the construction of $\Omega_{m}$ and $\Omega_{1-m}$. Let $g_{0}(z)=-\frac{1}{z}+\frac{1}{2z^{2}}+\frac{1}{2}$, by Lemma 6.1 we know $g_{\epsilon}(z)$ converges uniformly to $g_{0}(z)$ as $\epsilon\rightarrow 0$ on $|z|\geq\frac{1}{2}$. One can see $g_{0}(1)=0$, and $\forall\zeta>0$, $\exists C>0$, such that on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$, we have $g_{0}(x)>C$. Due to Lemma 6.1, we have $g_{\epsilon}(x)>\frac{C}{2}>0$ on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$ for $\epsilon$ small enough (i.e. for $m$ small enough since $\epsilon=\frac{R}{\eta}=\frac{Rm^{2}(1-m)^{2}}{J^{2}}$). Thus (3.3.1) implies $x=\frac{d}{\eta}$ has to satisfy $1-\zeta<\frac{d}{\eta}<1+\zeta$, which proves the proposition. ∎